# Double robust estimation of functional outcomes with data missing at random

### v2

@@ 3 Monte Carlo study
We here present numerical experiments to study finite sample properties of the estimators in relation to the asymptotic results given above. In particular, we study empirical coverages of the simultaneous confidence bands using the asymptotic covariance functions deduced in Theorems 2.3 and 2.4.
@@ 3.1 Design
We generate outcomes $\textbf{y}_{i}$ following (A.1) (in Appendix A) at $T=50$ equidistant points on $[0,1],$ using the following data generating process: $x_{1i}=1$; $(x_{2i},x_{3i},x_{4i})\sim\mathcal{MVN}(\mu,\Sigma)$ with $\mu=(-2,4,0)$,
@@
$x_{5i}\sim Ber(0.2)$; $x_{6i}\sim Bin(3,0.6)$. The functional coefficients in $\boldsymbol{\beta}$ are set to $40-t,2\sin(4t),3-\cos(5t),1.5\log(5t+0.1),0.5\sin(2t),2-1.5t+1.3t^{2}$, respectively, and evaluated at time $t_{j},j=1,\dots,T.$ The random error term is generated in the first situation as $\boldsymbol{\epsilon}_{i}\sim\mathcal{MVN}(\mathbf{0},\Sigma_{\epsilon})$. Here, $\Sigma_{\epsilon}$ is the discretisation of a (stationary) Matérn covariance function so that
@@
with $\kappa=1.5$, $\phi=0.1$, and where $K_{\kappa}$ is the modified Bessel function of the third kind of order $\kappa$. In a second situation, we generate the error term as $\boldsymbol{\epsilon}_{i}\sim\mathcal{MVT}(\mathbf{0},\Delta,\upsilon)$, so that the covariance matrix $\Sigma_{\epsilon}=\Delta\upsilon/(\upsilon-2)$, with $\upsilon=4$. The dispersion matrix $\Delta$ is generated as $Q\Lambda Q^{\top}$, where $Q$ is a $T\times T$ arbitrary orthonormal matrix and $\Lambda$ is a $T\times T$ diagonal matrix whose diagonal elements are a sequence of length $T$ ranging from 1 to 3. Finally, the missing indicator variable $Z_{i}$ is generated using (2), with $\boldsymbol{\gamma}=(0.3,-0.3,-0.3,-0.3,-0.3,-0.3)^{\top}$, so that $\Pr[Z_{i}=1|\textbf{x}_{i}]=0.71$. This means that on average, around 30% of the simulated functional outcomes are missing.
@@
We use 1000 replicates with sample sizes $n=250,500,1000$ and $3000$. For each replicate, we estimate $\mu_{y}$ with the outcome regression and double robust estimators. In order to confirm the double robustness property of the latter estimator, we consider four different scenarios. Firstly, when both working models, (A.1) and (2), are correctly specified; secondly, when only the outcome regression model, (A.1), is misspecified. Thirdly, when only the model for the missingness mechanism, (2), is misspecified, and lastly when both working models are misspecified. In each case, the misspecification entails that the model(s) omit(s) two of the covariates, $x_{3}$ and $x_{5}$. The performance of the estimators and their inference is evaluated with respect to bias, variance, mean squared error (MSE), and the empirical coverage of 95% confidence bands, both pointwise (PCB) and simultaneous bands (SCB). The SCB is constructed using sample variants of the covariance functions given in Theorem 2.3 and 2.4. This is done using the R-package ffscb (Liebl and Reimherr, 2019), implementing the computation of the functional critical value $\mathit{u}^{\star}_{.025}(t)$. The PCB instead uses a constant critical value $t_{.025}$, the 97.5th percentile from the t-distribution with n1n-1 degrees of freedom. In addition, we compare the performance of the estimators to the naive estimate based on complete case analysis, i.e. calculating the sample mean and its covariance based only the observed functional outcomes.
@@ 3.2 Results
Figure C.1 in Appendix C illustrates the width of the pointwise compared to the simultaneous confidence bands. Figures C.3-C.7 in the Appendix C display biases, variances, and MSE for the estimators in the different scenarios. There we can see that, as expected, the bias of the DR estimator is negligible when one or both of the two working models are correctly specified, similarly for the OR estimator when the outcome model is correct. When both models are misspecified, the DR estimator displays some bias, which varies between 0.17 and 0.38. A similar bias is observed for the OR estimator when the outcome model is incorrect. For comparison, the bias of the naive estimate varies between 0.2 and 1.39. These results confirm the expected double robustness properties of the DR estimator.
@@ 4 Application: Answering a counterfactual question
To illustrate the introduced methods we revisit a cohort dataset including all people born in Sweden year 1954. This data was used earlier in Ecker et al. (2024) to contrast the lifetime income trajectories between those living at 20 years of age in large labour market regions compared with those living in small labour market regions. Local labour markets are units defined by patterns of employment and commuting at the municipality level (Karlsson and Olsson, 2006; Statistics Sweden, 2010).
@@
More precisely, the outcome of interest $\cal Y$ measured annually is the logarithm of the accumulated total earned income in SEK, defined as all taxable income (except capital). We adjust incomes for inflation to match the monetary value in 2017, discounting with a factor of 0.03 (Johansson and Kriström, 2016).
@@
We estimate $\mu_{y}$ with the double robust estimator. In Figure 2 we display the comparison $\hat{\mu}_{DR}-\widehat{E}({\cal Y}\mid Z=1)$, i.e. comparing $\hat{\mu}_{DR}$ to the life income trajectory of those that actually started their labour market career in one of the large city. We see with this comparison that the naive estimation overestimates the income trajectory in the early years, but already before 30 years of age, it starts underestimating accumulated incomes. Confidence bands are obtained using the results of Theorem 2.4 and (8) applied to the double robust estimator.
@@ 5 Discussion
In this paper, we have introduced and studied two semi-parametric estimators for the mean of functional outcomes in the presence of missing (at random) outcome data. We have shown that both of these estimators (an outcome regression- and a double robust estimator) are asymptotically Gaussian processes, and presented expressions for their covariance functions. This enables us to quantify the uncertainty of these estimators by constructing simultaneous confidence bands with asymptotically guaranteed coverage. These asymptotics translate into satisfactory finite sample performances in the Monte Carlo study conducted. Further, the estimators studied here can be used in the context of causal inference from observational data, using the potential outcomes framework (Ding and Li, 2018), as illustrated by the application in Section 4.
A central assumption we have made is that the functional outcomes are missing at random. While this assumption is commonly required in the literature, it is nevertheless restrictive and cannot be tested without further assumptions (Molenberghs et al., 2015). Future work could thus consider developing uncertainty intervals that take into account departures from the missing at random assumption (Vansteelandt et al., 2006; Genbäck et al., 2015). Other possible extensions include considering the targeted learning estimator (van der Laan and Rose, 2018), which is closely related to the double robust estimator, and more generally, developing semi-parametric theory in the context of functional outcomes so that a larger class of machine learning methods can be used to fit the nuisance models.
{acks}
[Acknowledgments] We are thankful for comments by David Källberg that have improved the paper.
{funding}
This work was funded by a grant from the Marianne and Marcus Wallenberg Foundation.
@@ Appendix A Asymptotic results for finite-dimensional distributions
$\textbf{y}_{i}=\textbf{B}^{\top}\textbf{x}_{i}+\boldsymbol{\epsilon}_{i}$ (A.1)
where $\textbf{B}=(\textbf{b}_{1},\dots,\textbf{b}_{T})$, $\textbf{b}_{j}$ is a $p\times 1$ vector, and $\boldsymbol{\epsilon}_{i}$ is a random vector with zero mean vectors and $\boldsymbol{\Sigma}_{\epsilon}$ as the covariance matrix. The OR estimator of $\boldsymbol{\mu}_{y}=\mathbb{E}(\textbf{y})$ based on $(\textbf{y}_{i},\textbf{x}_{i},Z_{i})_{i=1}^{n}$ and (A.1) is
$\hat{\boldsymbol{\mu}}_{OR}=n^{-1}\sum_{i=1}^{n}\hat{\textbf{B}}^{\top}\textbf{x}_{i},$ (A.2)
where $\hat{\textbf{B}}$ is the ordinary least squares estimator. Similarly, based on the two working models (2) and (A.1), the corresponding double robust estimator is
$\hat{\boldsymbol{\mu}}_{DR}=n^{-1}\sum_{i=1}^{n}\left(\hat{\textbf{B}}^{\top}\textbf{x}_{i}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\boldsymbol{\gamma}})}(\textbf{y}_{i}-\hat{\textbf{B}}^{\top}\textbf{x}_{i})\right),$ (A.3)
where $\hat{\boldsymbol{\gamma}}$ again is the maximum likelihood estimator. The above estimators can be viewed as M-estimators and we assume throughout the usual regularity conditions for such estimators to be well-behaved asymptotically; see Newey and McFadden (1994) and Cantoni and de Luna (2020, Suppl Material).
Lemma A.1.
Assume that at least one of the two working models (1) and (2) is correctly specified, i.e., $E(\textbf{y}_{i}\mid\textbf{x}_{i})=\textbf{B}^{\top}\textbf{x}_{i}$ or $\Pr[Z_{i}=1|\textbf{x}_{i}]=\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})$. Then, $\sqrt{n}(\hat{\boldsymbol{\mu}}_{DR}-\boldsymbol{\mu}_{y})$ has an asymptotic multivariate normal distribution with mean zero. In the case that both working models are correctly specified, the covariance matrix simplifies and we have:
$\sqrt{n}(\hat{\boldsymbol{\mu}}_{DR}-\boldsymbol{\mu}_{y})\overset{d}{\rightarrow}\mathcal{N}\left(\textbf{0},\textbf{B}^{\top}\boldsymbol{\Sigma}_{x}\textbf{B}+\mathbb{E}[\tau^{-1}(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})]\boldsymbol{\Sigma}_{\epsilon}\right).$
@@
Based on the two models, the double robust estimator of the expected value of the outcome variables, $\boldsymbol{\mu}_{y}=\mathbb{E}(\textbf{y})$, is defined as
$\hat{\boldsymbol{\mu}}_{DR}=n^{-1}\sum_{i=1}^{n}\left(\hat{\textbf{B}}^{\top}\textbf{x}_{i}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\boldsymbol{\gamma}})}(\textbf{y}_{i}-\hat{\textbf{B}}^{\top}\textbf{x}_{i})\right),$
@@
where $\hat{\textbf{B}}$ and $\hat{\boldsymbol{\gamma}}$ are maximum likelihood estimators. We apply the M-estimation approach to derive the asymptotic distribution of the double robust estimator. To do so, we first define the vector equations $G_{n}(\boldsymbol{\theta})=n^{-1}\sum_{i=1}^{n}{\psi_{i}(D_{i},\boldsymbol{\theta})}$, where $D_{i}$ denotes the data of $i$th observation, and $\boldsymbol{\theta}=(\boldsymbol{\mu}_{y},\emph{vec}(\textbf{B}),\boldsymbol{\gamma})^{\top}$ denotes the vector of all parameters. To simplify the notations, we ignore $D_{i}$ in the equation $\psi_{i}$. The estimator $\hat{\boldsymbol{\theta}}$, which is the solution of the vector equations $G_{n}(\boldsymbol{\theta})=0$, has the following limiting distribution:
$\sqrt{n}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0})\to\mathcal{MN}(\textbf{0},\boldsymbol{\Sigma}_{\theta}),$
@@
where $\boldsymbol{\Sigma}_{\theta}=\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)^{-1}\mathbb{E}[\psi_{i}(\boldsymbol{\theta}_{0})\psi_{i}(\boldsymbol{\theta}_{0})^{\top}]\left\{\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)^{-1}\right\}^{\top}$.
@@
In order to define the estimating equations in terms of $\emph{vec}(\textbf{B})$, we need to introduce an equivalent representation of the model (A.1), which is:
$\textbf{y}_{i}=(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B})+\boldsymbol{\epsilon}_{i}.$
@@
Based on this representation we can define the following equations, which together make up $\psi_{i}(\boldsymbol{\theta})$:
$\psi_{1,i}(\boldsymbol{\theta})=(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B})+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})}(\textbf{y}_{i}-(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}))-\boldsymbol{\mu}_{y}$ (A.4)
$\psi_{2,i}(\boldsymbol{\theta})=(\textbf{y}_{i}-(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}))\otimes Z_{i}\textbf{x}_{i}$ (A.5)
$\psi_{3,i}(\boldsymbol{\theta})=\frac{Z_{i}-\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})(1-\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}))}\frac{\partial\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})}{\partial\boldsymbol{\gamma}}.$ (A.6)
@@
Note: (A.4), (A.5) and (A.6) all are vector valued functions and the dimensions are $T\times 1$, $Tp\times 1$, and $p\times 1$, respectively. Calculate the Jacobian of $\psi_{i}(\boldsymbol{\theta})$ and evaluate it at the true value $\boldsymbol{\theta}_{0}$. The first block of rows of the Jacobian is
@@
$\frac{\partial\psi_{1,i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}^{\top}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}=\left[-\textbf{I}_{T\times T},\frac{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})-Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top}),-Z_{i}e^{-\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}}\boldsymbol{\epsilon}_{i}\textbf{x}_{i}^{\top}\right].$
@@
$\frac{\partial\psi_{2,i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}^{\top}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}=\left[\textbf{0}_{Tp\times T},Z_{i}\textbf{I}_{T\times T}\otimes\textbf{x}_{i}\textbf{x}_{i}^{\top},\textbf{0}_{p\times p}\right].$
@@
$\frac{\partial\psi_{3,i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}^{\top}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}=\left[\textbf{0}_{p\times T},\textbf{0}_{p\times Tp},\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})(1-\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}))\textbf{x}_{i}\textbf{x}_{i}^{\top}\right].$

### v3

@@ 3 Monte Carlo study
We here present numerical experiments to study finite sample properties of the estimators in relation to the asymptotic results given above. In particular, we study empirical coverages of the simultaneous confidence bands using the asymptotic covariance functions deduced in Theorems 1 and 2.
@@ 3.1 Design
We generate outcomes $\textbf{y}_{i}$ following (9) (in Appendix A) at $T=50$ equidistant points on $[0,1],$ using the following data generating process: $x_{1i}=1$; $(x_{2i},x_{3i},x_{4i})\sim\mathcal{MVN}(\mu,\Sigma)$ with $\mu=(-2,4,0)$,
@@
$x_{5i}\sim Ber(0.2)$; $x_{6i}\sim Bin(3,0.6)$. The functional coefficients in $\bm{\beta}$ are set to $40-t,2\sin(4t),3-\cos(5t),1.5\log(5t+0.1),0.5\sin(2t),2-1.5t+1.3t^{2}$, respectively, and evaluated at time $t_{j},j=1,\dots,T.$ The random error term is generated in the first situation as $\bm{\epsilon}_{i}\sim\mathcal{MVN}(\mathbf{0},\Sigma_{\epsilon})$. Here, $\Sigma_{\epsilon}$ is the discretisation of a (stationary) Matérn covariance function so that
@@
with $\kappa=1.5$, $\phi=0.1$, and where $K_{\kappa}$ is the modified Bessel function of the third kind of order $\kappa$. In a second situation, we generate the error term as $\bm{\epsilon}_{i}\sim\mathcal{MVT}(\mathbf{0},\Delta,\upsilon)$, so that the covariance matrix $\Sigma_{\epsilon}=\Delta\upsilon/(\upsilon-2)$, with $\upsilon=4$. The dispersion matrix $\Delta$ is generated as $Q\Lambda Q^{\top}$, where $Q$ is a $T\times T$ arbitrary orthonormal matrix and $\Lambda$ is a $T\times T$ diagonal matrix whose diagonal elements are a sequence of length $T$ ranging from 1 to 3. Finally, the missing indicator variable $Z_{i}$ is generated using (2), with $\bm{\gamma}=(0.3,-0.3,-0.3,-0.3,-0.3,-0.3)^{\top}$, so that $\Pr[Z_{i}=1]=0.71$. This means that on average, around 30% of the simulated functional outcomes are missing.
@@
We use 1000 replicates with sample sizes $n=250,500,1000$ and $3000$. For each replicate, we estimate $\mu_{y}$ with the outcome regression and doubly robust estimators. In order to confirm the doubly robustness property of the latter estimator, we consider four different scenarios. Firstly, when both working models, (9) and (2), are correctly specified; secondly, when only the outcome regression model, (9), is misspecified. Thirdly, when only the model for the missingness mechanism, (2), is misspecified, and lastly when both working models are misspecified. In each case, the misspecification entails that the model(s) omit(s) two of the covariates, $x_{3}$ and $x_{5}$. The performance of the estimators and their inference is evaluated with respect to bias, variance, mean squared error (MSE), and the empirical coverage of 95% confidence bands, both pointwise (PCB) and simultaneous bands (SCB). The SCB is constructed using sample variants of the covariance functions given in Theorem 1 and 2. This is done using the R-package ffscb ffscb , implementing the computation of the functional critical value $\mathit{u}^{\star}_{.025}(t)$. The PCB instead uses a constant critical value $t_{.025}$, the 97.5th percentile from the t-distribution with n1n-1 degrees of freedom. In addition, we compare the performance of the estimators to the naive estimate based on complete case analysis, i.e. calculating the sample mean and its covariance based only the observed functional outcomes.
@@ 3.2 Results
Figure 3 in Appendix C illustrates the width of the pointwise compared to the simultaneous confidence bands. Figures 5-9 in the Appendix C display biases, variances, and MSE for the estimators in the different scenarios. There we can see that, as expected, the bias of the DR estimator is negligible when one or both of the two working models are correctly specified, similarly for the OR estimator when the outcome model is correct. When both models are misspecified, the DR estimator displays some bias, which varies between 0.17 and 0.38. A similar bias is observed for the OR estimator when the outcome model is incorrect. For comparison, the bias of the naive estimate varies between 0.2 and 1.39. These results confirm the expected doubly robustness properties of the DR estimator.
@@ 4 Application: Answering a counterfactual question
To illustrate the introduced methods we revisit a cohort dataset including all people born in Sweden year 1954. This data was used earlier in ecker2023causal to contrast the lifetime income trajectories between those living at 20 years of age in large labour market regions compared with those living in small labour market regions. Local labour markets are units defined by patterns of employment and commuting at the municipality level LA2 ; LA .
@@
More precisely, the outcome of interest $\cal Y$ measured annually is the logarithm of the accumulated total earned income in SEK, defined as all taxable income (except capital). We adjust incomes for inflation to match the monetary value in 2017, discounting with a factor of 0.03 disk .
@@
We estimate $\mu_{y}$ with the doubly robust estimator. In Figure 2 we display the comparison $\hat{\mu}_{DR}-\widehat{E}({\cal Y}\mid Z=1)$, i.e. comparing $\hat{\mu}_{DR}$ to the life income trajectory of those that actually started their labour market career in one of the large city. We see with this comparison that the naive estimation overestimates the income trajectory in the early years, but already before 30 years of age, it starts underestimating accumulated incomes. Confidence bands are obtained using the results of Theorem 2 and (8) applied to the doubly robust estimator.
@@ 5 Discussion
In this paper, we have introduced and studied two semi-parametric estimators for the mean of functional outcomes in the presence of missing (at random) outcome data. We have shown that both of these estimators (an outcome regression- and a doubly robust estimator) are asymptotically Gaussian processes, and presented expressions for their covariance functions. This enables us to quantify the uncertainty of these estimators by constructing simultaneous confidence bands with asymptotically guaranteed coverage. These asymptotics translate into satisfactory finite sample performances in the Monte Carlo study conducted. Further, the estimators studied here can be used in the context of causal inference from observational data, using the potential outcomes framework missingcausal , as illustrated by the application in Section 4.
A central assumption we have made is that the functional outcomes are missing at random. While this assumption is commonly required in the literature, it is nevertheless restrictive and cannot be tested without further assumptions molenberghs2015handbook . Future work could thus consider developing uncertainty intervals that take into account departures from the missing at random assumption vansteelandt2006ignorance ; genback2015uncertainty . Other possible extensions include considering the targeted learning estimator (tmle, ), which is closely related to the doubly robust estimator, and more generally, developing semi-parametric theory in the context of functional outcomes so that a larger class of machine learning methods can be used to fit the nuisance models.
Acknowledgments
We are thankful for comments by David Källberg and two anonymous referees that have improved the paper. This work was funded by a grant from the Marianne and Marcus Wallenberg Foundation.
Appendix
@@ Appendix A Asymptotic results for finite-dimensional distributions
$\textbf{y}_{i}=\textbf{B}^{\top}\textbf{x}_{i}+\bm{\epsilon}_{i}$ (9)
where $\textbf{B}=(\textbf{b}_{1},\dots,\textbf{b}_{T})$, $\textbf{b}_{j}$ is a $p\times 1$ vector, and $\bm{\epsilon}_{i}$ is a random vector with zero mean vectors and $\bm{\Sigma}_{\epsilon}$ as the covariance matrix. The OR estimator of $\bm{\mu}_{y}=\mathbb{E}(\textbf{y})$ based on $(\textbf{y}_{i},\textbf{x}_{i},Z_{i})_{i=1}^{n}$ and (9) is
$\hat{\bm{\mu}}_{OR}=n^{-1}\sum_{i=1}^{n}\hat{\textbf{B}}^{\top}\textbf{x}_{i},$ (10)
where $\hat{\textbf{B}}$ is the ordinary least squares estimator. Similarly, based on the two working models (2) and (9), the corresponding doubly robust estimator is
$\hat{\bm{\mu}}_{DR}=n^{-1}\sum_{i=1}^{n}\left(\hat{\textbf{B}}^{\top}\textbf{x}_{i}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\bm{\gamma}})}(\textbf{y}_{i}-\hat{\textbf{B}}^{\top}\textbf{x}_{i})\right),$ (11)
where $\hat{\bm{\gamma}}$ again is the maximum likelihood estimator. The above estimators can be viewed as M-estimators and we assume throughout the usual regularity conditions for such estimators to be well-behaved asymptotically; see [25] and CANTONI2020108 [5, Suppl Material].
Lemma 3.
Assume that at least one of the two working models (1) and (2) is correctly specified, i.e., $E(\textbf{y}_{i}\mid\textbf{x}_{i})=\textbf{B}^{\top}\textbf{x}_{i}$ or $\Pr[Z_{i}=1|\textbf{x}_{i}]=\tau(\textbf{x}_{i}^{\top}\bm{\gamma})$. Then, $\sqrt{n}(\hat{\bm{\mu}}_{DR}-\bm{\mu}_{y})$ has an asymptotic multivariate normal distribution with mean zero. In the case that both working models are correctly specified, the covariance matrix simplifies and we have:
$\sqrt{n}(\hat{\bm{\mu}}_{DR}-\bm{\mu}_{y})\overset{d}{\rightarrow}\mathcal{N}\left(\textbf{0},\textbf{B}^{\top}\bm{\Sigma}_{x}\textbf{B}+\mathbb{E}[\tau^{-1}(\textbf{x}_{i}^{\top}\bm{\gamma})]\bm{\Sigma}_{\epsilon}\right).$
@@
Based on the two models, the doubly robust estimator of the expected value of the outcome variables, $\bm{\mu}_{y}=\mathbb{E}(\textbf{y})$, is defined as
$\hat{\bm{\mu}}_{DR}=n^{-1}\sum_{i=1}^{n}\left(\hat{\textbf{B}}^{\top}\textbf{x}_{i}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\hat{\bm{\gamma}})}(\textbf{y}_{i}-\hat{\textbf{B}}^{\top}\textbf{x}_{i})\right),$
@@
where $\hat{\textbf{B}}$ and $\hat{\bm{\gamma}}$ are maximum likelihood estimators. We apply the M-estimation approach to derive the asymptotic distribution of the doubly robust estimator. To do so, we first define the vector equations $G_{n}(\bm{\theta})=n^{-1}\sum_{i=1}^{n}{\psi_{i}(D_{i},\bm{\theta})}$, where $D_{i}$ denotes the data of $i$th observation, and $\bm{\theta}=(\bm{\mu}_{y},\emph{vec}(\textbf{B}),\bm{\gamma})^{\top}$ denotes the vector of all parameters. To simplify the notations, we ignore $D_{i}$ in the equation $\psi_{i}$. The estimator $\hat{\bm{\theta}}$, which is the solution of the vector equations $G_{n}(\bm{\theta})=0$, has the following limiting distribution:
$\sqrt{n}(\hat{\bm{\theta}}-\bm{\theta}_{0})\to\mathcal{MN}(\textbf{0},\bm{\Sigma}_{\theta}),$
@@
where $\bm{\Sigma}_{\theta}=\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)^{-1}\mathbb{E}[\psi_{i}(\bm{\theta}_{0})\psi_{i}(\bm{\theta}_{0})^{\top}]\left\{\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)^{-1}\right\}^{\top}$.
@@
In order to define the estimating equations in terms of $\emph{vec}(\textbf{B})$, we need to introduce an equivalent representation of the model (9), which is:
$\textbf{y}_{i}=(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B})+\bm{\epsilon}_{i}.$
@@
Based on this representation we can define the following equations, which together make up $\psi_{i}(\bm{\theta})$:
$\psi_{1,i}(\bm{\theta})=(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B})+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma})}(\textbf{y}_{i}-(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}))-\bm{\mu}_{y}$ (12)
$\psi_{2,i}(\bm{\theta})=(\textbf{y}_{i}-(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}))\otimes Z_{i}\textbf{x}_{i}$ (13)
$\psi_{3,i}(\bm{\theta})=\frac{Z_{i}-\tau(\textbf{x}_{i}^{\top}\bm{\gamma})}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma})(1-\tau(\textbf{x}_{i}^{\top}\bm{\gamma}))}\frac{\partial\tau(\textbf{x}_{i}^{\top}\bm{\gamma})}{\partial\bm{\gamma}}.$ (14)
@@
Note: (12), (13) and (14) all are vector valued functions and the dimensions are $T\times 1$, $Tp\times 1$, and $p\times 1$, respectively. Calculate the Jacobian of $\psi_{i}(\bm{\theta})$ and evaluate it at the true value $\bm{\theta}_{0}$. The first block of rows of the Jacobian is
@@
$\frac{\partial\psi_{1,i}(\bm{\theta})}{\partial\bm{\theta}^{\top}}\biggr\rvert_{\bm{\theta}_{0}}=\left[-\textbf{I}_{T\times T},\frac{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})-Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top}),-Z_{i}e^{-\textbf{x}_{i}^{\top}\bm{\gamma}_{0}}\bm{\epsilon}_{i}\textbf{x}_{i}^{\top}\right].$
@@
$\frac{\partial\psi_{2,i}(\bm{\theta})}{\partial\bm{\theta}^{\top}}\biggr\rvert_{\bm{\theta}_{0}}=\left[\textbf{0}_{Tp\times T},Z_{i}\textbf{I}_{T\times T}\otimes\textbf{x}_{i}\textbf{x}_{i}^{\top},\textbf{0}_{p\times p}\right].$
@@
$\frac{\partial\psi_{3,i}(\bm{\theta})}{\partial\bm{\theta}^{\top}}\biggr\rvert_{\bm{\theta}_{0}}=\left[\textbf{0}_{p\times T},\textbf{0}_{p\times Tp},\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})(1-\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0}))\textbf{x}_{i}\textbf{x}_{i}^{\top}\right].$